# Online, Target-Free LiDAR–Camera Extrinsic Calibration via Cross-Modal Mask Matching

### v1

@@ IV-B Datasets
We have created the following three real-world datasets: MIAS-LCEC-TL70 (target-free), MIAS-LCEC-CB70 (target-based), and MIAS-LCEC-TL360 (target-free), which are now publicly available for researchers to evaluate the performance of LCEC approaches:
MIAS-LCEC-TL70 is a diverse and challenging dataset that contains 60 pairs of 4D point clouds (including spatial coordinates with intensity data) and RGB images, collected using a Livox Mid-70 LiDAR and a MindVision SUA202GC camera, from a variety of indoor and outdoor environments, under various scenarios as well as different weather and illumination conditions. We divide this dataset into six subsets: residential community, urban freeway, building, challenging weather, indoor, and challenging illumination, to comprehensively evaluate the algorithm performance.
@@
MIAS-LCEC-TL360 contains 10 pairs of 4D point clouds and RGB images, collected using a Livox Mid-360 LiDAR and a MindVision SUA202GC camera from both indoor and outdoor environments. Since the Livox Mid-360 LiDAR has a scanning range of 360∘, it produces a sparser point cloud compared to that generated by the Livox Mid-70 LiDAR. Additionally, the significant difference in the FoV between this type of LiDAR and the camera results in only a small overlap in the collected data. Consequently, this dataset is particularly well-suited for evaluating the adaptability of algorithms to challenging scenarios characterized by sparse point clouds and limited data overlap.
@@ IV-C Evaluation Metrics
between 3D LiDAR points and their corresponding 2D image pixels to quantify the LCEC algorithms when using targets.
@@
(∘) | HKU-Mars [14] | 74.719 | 67.558 |
DVL [17] | 60.041 | 59.729 | |
MIAS-LCEC (Ours) | 1.042 | 1.931 | |
(m) | HKU-Mars [14] | 2.711 | 6.833 |
DVL [17] | 0.980 | 3.127 | |
MIAS-LCEC (Ours) | 0.207 | 0.236 |
@@ IV-D Comparisons with State-of-the-Art Methods
Quantitative comparisons with SoTA approaches on the MIAS-LCEC-TL70 and MIAS-LCEC-TL360 datasets are presented in Tables I and II. Additionally, qualitative results for these datasets are illustrated in Figs. 5, 6, and 7. It is important to note that the results from the first iteration of MIAS-LCEC are reported here because the accuracy achieved is already higher than that of existing SoTA approaches.
The results shown in Table I suggest that our method outperforms both HKU-Mars and DVL across a total of 60 scenarios, all captured using a Livox Mid-70 LiDAR. Specifically, MIAS-LCEC reduces by around 22-88% and decreases by 40-95%, compared to existing SoTA algorithms. We attribute these performance improvements to the coarse-to-fine correspondence matching pipeline based on LVMs, which sets strict criteria for reliable sparse correspondence selection and propagates these matches to generate dense correspondences, thereby improving the quality of the PnP solutions. It can also be observed that MIAS-LCEC achieves lower mean values than both HKU-Mars and DVL across all six subsets. While the mean achieved by our method is slightly higher (around ) than that of DVL in the indoor subset, MIAS-LCEC outperforms both algorithms in terms of mean in the remaining five subsets. Our method dramatically outperforms HKU-Mars and is slightly better than DVL in scenarios with low noise and abundant features, while it performs significantly better than both methods in challenging conditions, particularly under poor illumination and adverse weather, or when few geometric features are detectable. This impressive performance can be attributed to MobileSAM, a powerful LVM, capable of learning informative, general-purpose deep features for robust image segmentation. Surprisingly, as observed in Fig. 5, MobileSAM can effectively segment both RGB and LIP images captured in challenging conditions, such as dark underground parking garages or during rainy days, where the objects are even unrecognizable to human observers.
@@
Offline [28] | 1.352 | 1.343 | 1.346 | 1.347 | |
MIAS-LCEC (Ours) | 1.369 | 1.484 | 1.640 | 1.498 | |
Offline [28] | 1.768 | 1.641 | 1.429 | 1.613 | |
MIAS-LCEC (Ours) | 2.023 | 1.832 | 1.474 | 1.776 | |
Offline [28] | 1.727 | 1.712 | 1.499 | 1.646 | |
MIAS-LCEC (Ours) | 1.740 | 1.886 | 1.740 | 1.789 | |
Offline [28] | 1.568 | 1.965 | 1.748 | 1.760 | |
MIAS-LCEC (Ours) | 1.579 | 1.830 | 2.168 | 1.859 | |
Offline [28] | 1.404 | 1.477 | 1.456 | 1.446 | |
MIAS-LCEC (Ours) | 1.491 | 2.025 | 1.889 | 1.802 |
Additionally, the results on the MIAS-LCEC-TL360 dataset somewhat exceed our expectations. From Table II, it is evident that while HKU-Mars and DVL achieve poor performance on this dataset, MIAS-LCEC demonstrates acceptable performance, indicating strong adaptability to more challenging scenarios, with narrow overlapping areas between LIP and RGB images. This performance improvement is primarily attributed to our developed LIP image generation strategy, which incorporates several image pre-processing techniques in our practical implementation to refine the LIP images and align the FoVs between the two sensors as closely as possible. As illustrated in Figs. 6 and 7, the LIP image generated by DVL contains numerous holes and has a significantly different FoV compared to the RGB image, resulting in unexpected false correspondence matches, which can deteriorate the algorithm’s efficiency and robustness. In contrast, MIAS-LCEC can generate LIP images that look as if taken from the same perspective to the actual camera, thus improving the performance of cross-modal mask matching.
IV-E Comparison with An Offline, Target-Based Approach
@@ IV-F LCEC Performance with Increasing Iterations
Fig. 9 illustrates the accuracy of our algorithm with respect to an increasing number of iterations. It is evident that (1) after the first iteration, our approach attains satisfactory accuracy, and (2) after the third iteration, its performance stabilizes and remains relatively consistent (the values of and decrease by approximately 16% and 11%, respectively, from the first to the sixth iterations). Therefore, we contend that a single iteration suffices for our MIAS-LCEC approach, striking a balance between accuracy and efficiency. However, additional iterations can certainly be considered when computational resources are abundant.

### v2

@@ IV-B Datasets
We have created the following three real-world datasets: MIAS-LCEC-TF70 (target-free), MIAS-LCEC-CB70 (target-based), and MIAS-LCEC-TF360 (target-free), which are now publicly available for researchers to evaluate the performance of LCEC approaches:
MIAS-LCEC-TF70 is a diverse and challenging dataset that contains 60 pairs of 4D point clouds (including spatial coordinates with intensity data) and RGB images, collected using a Livox Mid-70 LiDAR and a MindVision SUA202GC camera, from a variety of indoor and outdoor environments, under various scenarios as well as different weather and illumination conditions. We divide this dataset into six subsets: residential community, urban freeway, building, challenging weather, indoor, and challenging illumination, to comprehensively evaluate the algorithm performance.
@@
MIAS-LCEC-TF360 contains 12 pairs of 4D point clouds and RGB images, collected using a Livox Mid-360 LiDAR and a MindVision SUA202GC camera from both indoor and outdoor environments. Since the Livox Mid-360 LiDAR has a scanning range of 360∘, it produces a sparser point cloud compared to that generated by the Livox Mid-70 LiDAR. Additionally, the significant difference in the FoV between this type of LiDAR and the camera results in only a small overlap in the collected data. Consequently, this dataset is particularly well-suited for evaluating the adaptability of algorithms to challenging scenarios characterized by sparse point clouds and limited data overlap.
@@ IV-C Evaluation Metrics
between 3D LiDAR points and their corresponding 2D image pixels to quantify the performance of LCEC algorithms when using targets.
@@
(∘) | CRLF [25] | 1.469 | 1.402 |
UMich [43] | 1.802 | 2.698 | |
HKU-Mars [19] | 96.955 | 25.611 | |
DVL [22] | 63.003 | 46.623 | |
MIAS-LCEC (Ours) | 0.963 | 0.659 | |
(m) | CRLF [25] | 13.484 | 0.139 |
UMich [43] | 0.200 | 0.135 | |
HKU-Mars [19] | 4.382 | 9.914 | |
DVL [22] | 0.919 | 1.778 | |
MIAS-LCEC (Ours) | 0.182 | 0.114 |
@@ IV-D Comparisons with State-of-the-Art Methods
Quantitative comparisons with SoTA approaches on the MIAS-LCEC-TF70 and MIAS-LCEC-TF360 datasets are presented in Tables I and II. Additionally, qualitative results for these datasets are illustrated in Figs. 6, 7, and 8. It is important to note that the results from the first iteration of MIAS-LCEC are reported here because the accuracy achieved is already higher than that of existing SoTA approaches.
The results shown in Table I suggest that our method outperforms all other SoTA approaches on a total of 60 scenarios, all captured using a Livox Mid-70 LiDAR. Specifically, MIAS-LCEC reduces by around 30-93% and decreases by 39-99%, compared to existing SoTA algorithms. We attribute these performance improvements to the coarse-to-fine correspondence matching pipeline based on LVMs, which sets strict criteria for reliable sparse correspondence selection and propagates these matches to generate dense correspondences, thereby improving the quality of the PnP solutions. It can also be observed that MIAS-LCEC achieves lower mean and values than all other approaches across the total six subsets. Our method dramatically outperforms CRLF, UMich, and HKU-Mars and is slightly better than DVL in scenarios with low noise and abundant features, while it performs significantly better than all methods in challenging conditions, particularly under poor illumination and adverse weather, or when few geometric features are detectable. This impressive performance can be attributed to MobileSAM, a powerful LVM, capable of learning informative, general-purpose deep features for robust image segmentation. Surprisingly, as observed in Fig. 6, MobileSAM can effectively segment both RGB and LIP images captured in challenging conditions, such as dark underground parking garages or during rainy days, where the objects are even unrecognizable to human observers.
Additionally, the results on the MIAS-LCEC-TF360 dataset somewhat exceed our expectations. From Table II, it is evident that while the other approaches achieve poor performance on this dataset, MIAS-LCEC demonstrates acceptable performance, indicating strong adaptability to more challenging scenarios, with narrow overlapping areas between LIP and RGB images. This performance improvement is primarily attributed to our developed LIP image generation strategy, which incorporates several image pre-processing techniques in our practical implementation to refine the LIP images and align the FoVs between the two sensors as closely as possible. As illustrated in Figs. 7 and 8, the LIP image generated by DVL contains numerous holes and has a significantly different FoV compared to the RGB image, resulting in unexpected false correspondence matches, which can deteriorate the algorithm’s efficiency and robustness. In contrast, MIAS-LCEC can generate LIP images that look as if taken from the same perspective to the actual camera, thus improving the performance of cross-modal mask matching.
IV-E Comparison with An Offline, Target-Based Approach
@@
Offline [33] | 1.451 | 1.359 | 1.307 | 1.372 | |
MIAS-LCEC (Ours) | 1.445 | 1.626 | 1.653 | 1.575 | |
Offline [33] | 1.589 | 1.544 | 1.361 | 1.498 | |
MIAS-LCEC (Ours) | 1.919 | 1.473 | 1.386 | 1.593 | |
Offline [33] | 1.664 | 2.132 | 1.329 | 1.708 | |
MIAS-LCEC (Ours) | 1.593 | 2.086 | 1.608 | 1.762 | |
Offline [33] | 1.539 | 1.838 | 1.743 | 1.706 | |
MIAS-LCEC (Ours) | 1.572 | 1.802 | 2.204 | 1.859 | |
Offline [33] | 1.439 | 1.534 | 1.470 | 1.481 | |
MIAS-LCEC (Ours) | 1.569 | 2.020 | 1.883 | 1.824 |
@@ IV-F LCEC Performance with Increasing Iterations
Fig. 10 illustrates the accuracy of our algorithm with respect to an increasing number of iterations. It is evident that (1) after the first iteration, our approach attains satisfactory accuracy, and (2) after the third iteration, its performance stabilizes and remains relatively consistent (the values of and decrease by approximately 26% and 11%, respectively, from the first to the sixth iterations). Therefore, we contend that a single iteration suffices for our MIAS-LCEC approach, striking a balance between accuracy and efficiency. However, additional iterations can certainly be considered when computational resources are abundant.